The population of each classroom for face-to-face classes consists of the numbers 10, 11, 15, and 9. List all the sample size of 3 from this population and compute the mean of each sample.

1
Expert's answer
2022-05-24T16:12:55-0400

We have population values 9,10,11,15, population size N=4 and sample size n=3.

Mean of population "(\mu)" = "\dfrac{9+10+11+15}{4}=11.25"

Select a random sample of size 3 without replacement. We have a sample distribution of sample mean.

The number of possible samples which can be drawn without replacement is "^{N}C_n=^{4}C_3=4."

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 no & Sample & Sample \\
& & mean\ (\bar{x})
\\ \hline
 1 & 9,10,11 & 30/3 \\
 \hdashline
 2 & 9,10,15 & 34/3 \\
 \hdashline
 3 & 9,11,15 & 35/3 \\
 \hdashline
 4 & 10,11,15 & 36/3 \\
 \hdashline
\end{array}"





"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \bar{X} & f(\bar{X}) &\bar{X} f(\bar{X}) & \bar{X}^2f(\bar{X})
\\ \hline
 30/3 & 1/4 & 30/12 & 900/36 \\
 \hdashline
 34/3 & 1/4 & 34/12 & 1156/36 \\
 \hdashline
 35/3 & 1/4& 35/12 & 1225/36 \\
 \hdashline
 36/3 & 1/4 & 36/12 & 1296/36 \\
 \hdashline
\end{array}"



Mean of sampling distribution 



"\mu_{\bar{X}}=E(\bar{X})=\sum\bar{X}_if(\bar{X}_i)=\dfrac{135}{12}=11.25=\mu"